A solution of 54% fertilizer is to be mixed with a solution of 27% fertilizer to form 135 liters of a 49% solution. How many liters of the 54% solution must be used?
Answer by mananth(16946)

---------percent ---------------- quantity
Fertlizer I 54 ---------------- x liters
Fertilizer II 27 ------ 135 - x liters
Mixture 49.00% ---------------- 135

54 x + 27 ( 135 - x ) = 49 * 135

54 x + 3645 - 27 x = 6615
54 x - 27 x = 6615 - 3645
27 x = 2970
/ 27
x = 110 liters 54.00% Fertlizer I
25 liters 27.00% Fertilizer II
m.ananth@hotmail.ca